# FAVAE-Effective Frequency Aware Latent Tokenizer

### v1

@@ 3 Experiments
Our analysis in Table 1 shows that KL-regularized VAEs with carefully designed parameterizations consistently outperform both standard VAEs and VQ-VAE baselines across all reconstruction metrics, achieving efficient performance at higher latent compression rates due to the absence of quantization artifacts. Their effectiveness is strongly tied to latent parameterization quality, as seen in recent autoregressive models chen2025masked and VA-VAE yao2025vavae , which further align latent spaces with foundation models (e.g., DINOv2 oquab2023dinov2 ) for improved perceptual quality. However, such models jointly optimize frequency components, leading to trade-offs and suboptimal preservation of low and high-frequency details (Especially high frequency details). In contrast, our FA-VAE explicitly decouples low- and high-frequency bands, enabling specialized representation learning and more precise reconstructions of both global structures and fine details. FA-VAE nearly halves the reconstruction loss of the strongest baseline while achieving the best overall performance across both low and high frequency reconstructions, demonstrating the benefits of frequency-aware modeling. The consistent gap between low- and high-frequency reconstructions across models further underscores the general difficulty of capturing fine details effectively. The other metrics like LPIPS and rFID showcases the perceptual quality of reconstructions of FA-VAE compared with other baselines.
4 Conclusion
@@
In this work, we investigate frequency awareness in the latent embeddings of latent tokenizer models. We find that jointly optimizing low- and high-frequency components leads to a frequency bias favoring low-frequencies thereby degrading the reconstruction of fine-grained, high-frequency details and overall perceptual quality. To address this limitation, we introduce FA-VAE, a frequency-aware VAE framework that explicitly decouples low and high-frequency components using wavelet decomposition, processes them independently, and fuses them into a unified latent representation. FA-VAE achieves state-of-the-art performance across frequency-aware reconstruction metrics, demonstrating improved fidelity and perceptual quality of the learned latent embeddings.
Appendix A Architecture and Implementation Details
@@
The architecture of our Frequency-Aware Variational Au- toencoder (FA-VAE) consists of two independent encoder- decoder pairs, designed to separately process low- and high-frequency components of the input data. The encoder- decoder pair architecture is inspired by VAVAE yao2025vavae . We further follow a similar hyperparameter setup to rombach2022high ; yao2025vavae for implementing our latent reconstruction module. To support distributed training across multiple nodes, we scale the learning rate and global batch size to $1\times 10^{-4}$ and 256, respectively, following the MAR setup li2024autoregressive . We experiment with two $f16$ tokenizers for low and high frequency subbands: one trained without visual alignment VF loss for high frequencies and one with VF loss using DINOv2 oquab2023dinov2 for low frequencies. Here, $f$ represents the latent spatial resolution factor and $d$ the latent dimensionality. Following yao2025vavae , we set the VF loss hyperparameters to $m_{1}=0.5$, $m_{2}=0.25$, and $w_{\text{hyper}}=0.1$. Figure 3 shows the qualitative reconstructions of FA-VAE.
Figure 3: Qualitative reconstructions using FA-VAE on ImageNet $\times256$.

### v2

@@ Introduction
Our analysis in Table 1 reveals that KL-regularized VAE with well-designed parameterizations consistently outperform both traditional VAE and vector quantization (VQ-VAE) baselines across all reconstruction metrics. Also KL-regularized VAE show efficient reconstruction performance at much better compression rate with lower dimensionality. This improvement stems from the absence of quantization artifacts, which often degrade reconstruction quality in discrete latent spaces. Additionally, the effectiveness of KL-VAE is closely tied to the quality of their latent parameterization schemes. Notably, KL-VAE employed in recent autoregressive models (Chen et al. 2025), and VA-VAE (Yao, Yang, and Wang 2025) which further aligns its latent space with foundation models like DINOv2 (Oquab et al. 2023) achieve strong performance across low and high-frequency reconstruction, along with improved perceptual quality. However, these models still rely on a joint optimization of frequency components, which can lead to trade-offs and suboptimal detail preservation, especially in the high-frequency spectrum. In contrast, our proposed Frequency-Aware VAE (FA-VAE) builds on the VA-VAE framework by explicitly decoupling and modeling low- and high-frequency components in learning of the compact latent embeddings. This separation enables focused representation learning for each frequency band, leading to more precise reconstructions of both global structures and fine details. FA-VAE achieves the best overall performance across all evaluated metrics nearly halving the reconstruction loss of the strongest baseline (VA-VAE) and highlights the benefit of frequency-aware modeling for learning more expressive, disentangled, and visually faithful latent embeddings. Notably, the performance gap between low and high-frequency reconstructions across all models highlights the general challenge of capturing fine details effectively.
Latent Generation Results
We evaluate the generative quality of our learned latent embeddings by training a diffusion-based generative model on top of the FA-VAE representations. Specifically, we use LightningDiT (Yao, Yang, and Wang 2025), a fast-converging latent diffusion backbone to model the generation process, and compare against a wide range of state-of-the-art autoregressive and latent diffusion models, both with and without classifier-free guidance (CFG). Table 2 reports quantitative results across standard metrics: generation FID score(gFID), structural FID (sFID), Inception Score (IS), Precision, and Recall. Lower gFID and sFID values indicate better image fidelity and diversity, while higher IS, precision, and recall reflect better sample quality.
Table 2 shows comparison of LightningDiT with FA-VAE with other standard latent generation baselines. Our approach, LightningDiT with FA-VAE, achieves the best gFID among all methods, along with highly competitive sFID and strong scores across IS, precision, and recall with and without the CFG setting. Compared to auto-regressive models like MaskGIT (Chang et al. 2022), MAR (Li et al. 2024a), VAR (Tian et al. 2024), MagViT-v2 (Yu et al. 2023) xand LlamaGen (Sun et al. 2024), as well as latent diffusion models such as MaskDiT (Zheng et al. 2023), DiT (Peebles and Xie 2023), SiT (Ma et al. 2024), FasterDiT (Yao et al. 2024), MDF (Gao et al. 2023) and REPA (Yu et al. 2024b), our method consistently delivers a better fidelity and diversity. Notably, it outperforms the LightningDiT baseline trained on VA-VAE embeddings across all metrics, demonstrating the advantage of our frequency-aware latent space in enhancing generation quality. These results validate the effectiveness of our FA-VAE framework for learning high-quality, structure-preserving representations that generalize well in generative pipelines. The qualitative results of our generated visualizations are shown in Figure 4.
Figure 5: Top-100 classes by reconstruction error. Our model shows consistently lower MSE across challenging categories.
Fairness in Latent Representations.
We evaluate fairness in terms of how well the latent embeddings represent the complex categories within the dataset. An embedding space is considered fair if it enables faithful reconstructions across all data categories, particularly those that are structurally complex. Since these embeddings are used in the generation pipeline, under-representation of certain classes can result in biased or degraded outputs. To analyze this, we identify the top 100 ImageNet categories with the highest reconstruction error, measured by normalized reconstruction error using mean squared error (NMSE). Figure 5 presents the class-wise reconstruction loss performance on these challenging categories. Our frequency-aware model consistently achieves lower reconstruction errors compared to VA-VAE, indicating improved fidelity and robustness in capturing complex details. These results demonstrate that frequency-aware modeling facilitates more balanced and faithful reconstructions, particularly in high-error regions of the data distribution.
Ablation Study
To further investigate the contribution of decoupling low- and high-frequency components in our frequency-aware training strategy, we conduct an ablation study utilizing few latent embedding tokenizers from Table 1. Specifically, we retrain these models by coupling the low- and high-frequency components in a shared frequency space, using the same wavelet-based losses employed in our FA-VAE framework. In this setup, each tokenizer is optimized jointly on both low- and high-frequency representations obtained via the Discrete Wavelet Transform (DWT) using Haar filters. The results, presented in Table 3, demonstrate that our frequency-aware FA-VAE consistently outperforms the coupled-frequency variants across all metrics, including overall reconstruction error (Rec.), low-(LF) and high-frequency (HF) fidelity error, perceptual similarity loss (LPIPS), and reconstruction FID (rFID). This indicates that explicitly decoupling frequencies during training leads to more accurate and perceptually faithful reconstructions. For a fair comparison, all models are trained with similar latent dimensionality as used in the FA-VAE configuration.
Conclusion
@@
In this work, we investigate frequency awareness in the latent embeddings of tokenizer models used within latent generative pipelines. We find that jointly optimizing low- and high-frequency components leads to a frequency bias favoring low-frequencies thereby degrading the reconstruction of fine-grained, high-frequency details and overall perceptual quality. To address this limitation, we introduce FA-VAE, a frequency-aware VAE framework that explicitly decouples low and high-frequency components via wavelet decomposition, processes them independently, and fuses them into a unified latent representation. FA-VAE achieves state-of-the-art performance across frequency-aware reconstruction metrics, demonstrating improved fidelity and perceptual quality.
Appendix A Appendix
In this appendix, we provide additional details supporting our main paper. The sections below elaborate on the architectural design, implementation specifics, and qualitative visualizations that complement our experimental findings:
  • Architectural details of FA-VAE
  • Implementation specifics and training setup
  • Visualization comparisons highlighting reconstruction quality
  • Additional Experiments
Appendix B Architectural Details of FA-VAE
The architecture of our Frequency-Aware Variational Autoencoder (FA-VAE) consists of two independent encoder-decoder pairs, designed to separately process low- and high-frequency components of the input data. The encoder-decoder architecture is inspired by standard VAE setups.
Shared Architectural Design
Each encoder and decoder follows a symmetric U-Net-style architecture with hierarchical multiscale processing. The core building blocks are ResnetBlocks with group normalization and Swish activations. Attention modules (AttnBlock or LinearAttention) are interleaved based on spatial resolution thresholds to capture long-range dependencies, particularly at intermediate scales such as $8\times 8$ or $16\times 16$. Downsampling in the encoder is performed via strided convolutional layers or average pooling, while upsampling in the decoder uses nearest-neighbor interpolation followed by convolution. The spatial resolution halves at each stage, yielding a compact bottleneck latent representation. This architectural design for individual encoder-decoder setup is inherited from  (Li et al. 2024a).
Low- and High-Frequency Encoding
To enable frequency-aware modeling, the input is first decomposed using the discrete wavelet transform (DWT) with a Haar filter at level 1 decomposition, yielding a low-frequency component (LL) and high-frequency components (LH, HL, HH). These frequency components are then processed separately:
  • Low-Frequency Encoder: This encoder receives the LL component and maps it through a standard hierarchical structure, outputting the parameters (mean and log-variance) of a diagonal Gaussian distribution. These parameters define the variational posterior from which low-frequency latent codes are sampled.
  • High-Frequency Encoder: The high-frequency tensor is reshaped and passed through a parallel encoder, similar in design to the low-frequency branch but adapted for higher input dimensionality (i.e., 9 channels for 3 scales $\times$ 3 orientations of LH, HL, HH). It also outputs Gaussian parameters and samples latent variables accordingly.
Both encoders use the MergedRescaleEncoder module, which applies additional residual blocks and attention after the initial latent encoding, allowing controlled compression and modulation of spatial structure.
Model Tokenizer Config Recon. Loss Low Freq. Loss High Freq. Loss LPIPS
DC-AE f32c32 0.0213 0.0521 0.0110 0.1708
DC-AE f32c32 0.0197 0.0478 0.0103 0.1704
DC-AE f64c128 0.0206 0.0499 0.0109 0.1689
DC-AE f128c512 0.0208 0.0510 0.0107 0.1756
KL-VAE f4c3 0.0064 0.0097 0.0053 0.0722
KL-VAE f8c4 0.0156 0.0348 0.0091 0.1459
KL-VAE f16c16 0.0163 0.0365 0.0096 0.1529
KL-VAE f32c64 0.0251 0.0640 0.0121 0.2089
KL-VAE f64c128 0.0241 0.0652 0.0105 0.1708
VQ-VAE f4c3v8192 0.0064 0.0104 0.0051 0.0693
VQ-VAE f8c4v16384 0.0203 0.0475 0.0113 0.1701
VQ-VAE f8c4v256 0.0239 0.0569 0.0128 0.1848
VQ-VAE f16c8v16384 0.0352 0.1006 0.0134 0.2637
KL-VAE f8c4 (ema) 0.0145 0.0313 0.0089 0.1345
[…]
Table 4: Comparison of tokenizers on the ImageNet validation set across frequency-aware and perceptual metrics. Lower values indicate better performance. f denotes the latent spatial downsampling factor (e.g., f=16f{=}16 implies $16\times$ downsampling in width and height), and c denotes the number of latent channels. For VQ models, v indicates the vocabulary size. Models trained on ImageNet, SAM, FFHQ, and Mapillary Vistas. Models trained on OpenImages. Models fine-tuned on OpenImages and LAION.
Bottleneck and Latent Space
Each encoder produces a bottleneck representation at a spatial scale determined by the number of downsampling layers (typically $16\times 16$ or $8\times 8$), with a corresponding increase in channel dimensionality. The encoder outputs are passed through $\times1$ convolutional layers to double the channels and parameterize a Gaussian distribution, following the VAE formulation (Van Den Oord, Vinyals et al. 2017). Sampling is performed using the reparameterization trick.
Decoding and Reconstruction
The sampled latent codes are passed through their respective decoders, which reverse the hierarchy of the encoder. The decoder structure mirrors the encoder in terms of resolution levels and attention placement. The outputs from the low- and high-frequency decoders are projected back to image space using a wavelet inversion step using inverse DWT.
Hierarchical Extensions
To improve semantic alignment and generalization, FA-VAE, inspired from VA-VAE (Yao, Yang, and Wang 2025) integrates the FirstStagePostProcessor, which enables conditioning on external features from a pretrained model (Dinov2). These modules enhance the latent space with multiscale guidance. This module is only active for low frequency encoded latents. During Generation, we model the concatenated low and high frequency latent embeddings using LightningDit (Yao, Yang, and Wang 2025). There are no architectural changes during the generation pipeline and only minimal adjustments in the auto-encoding pipeline.
Appendix C Implementation Details
@@
We follow a similar hyperparameter setup to (Rombach et al. 2022; Yao, Yang, and Wang 2025) for implementing our latent reconstruction and generation pipeline. For the visual tokenizer, we adopt the architectural backbone and training scheme from LDM (Rombach et al. 2022; Li et al. 2024a), leveraging a VQGAN-style encoder-decoder structure but omitting quantization to obtain a continuous latent space regulated by KL divergence. We apply the KL loss directly over the latent representation as in (Yao, Yang, and Wang 2025). To support distributed training across multiple nodes, we scale the learning rate and global batch size to $1\times 10^{-4}$ and 256, respectively, following the MAR setup (Li et al. 2024a). We experiment with two different $f16$ tokenizers: one trained without alignment for high frequencies and one with VF loss using DINOv2 (Oquab et al. 2023) for low frequencies. Here, $f$ represents the spatial resolution factor and $d$ the latent feature dimensionality. Following (Yao, Yang, and Wang 2025), we set the VF loss hyperparameters to $m_{1}=0.5$, $m_{2}=0.25$, and $w_{\text{hyper}}=0.1$.
For generative modeling, we employ LightningDiT (Yao, Yang, and Wang 2025), incorporating architectural enhancements and training strategies outlined in their work. The latent features extracted from the FA-VAE tokenizer are used to train LightningDiT models on ImageNet at $256\times 256$ resolution for 64 epochs. The patch size is set to 1 to ensure the effective downsampling rate of the entire system remains at 16, aligning with the approach in (Chen et al. 2024). We trained our models using 4 H100 Nvidia GPUs with 93.58GB VRAM.
Explanation of Metrics. We report several standard metrics commonly used to evaluate the quality of generated images from autoregressive and latent diffusion models.
  • gFID (generation FID): Frechet Inception Distance computed between generated images and real images, using the Inception-v3 network. Lower values indicate better alignment in feature distributions and overall visual fidelity.
  • sFID (spatial FID): A structural or spatial-aware variant of FID. sFID is equivalent to FID but uses intermediate spatial features in the inception network rather than the spatially-pooled features used in standard FID.
  • IS (Inception Score): Measures both the confidence and diversity of generated images based on class predictions from the Inception network. Higher IS indicates better generative diversity and visual sharpness.
  • Precision (Pre.) and Recall (Rec.): These metrics assess the coverage and fidelity of the generated data manifold with respect to real images. Higher precision implies fewer unrealistic samples (fidelity), while higher recall indicates better diversity and coverage.
All metrics are computed using 50K generated samples. We report results both with and without Classifier-Free Guidance (CFG) when available, which is a commonly used sampling technique in diffusion models to improve sample quality. Missing entries (marked as -) reflect the unavailability of results in the respective original papers.
Appendix D Visualization Comparisons
We qualitatively visualize the reconstructions of the best three models from Table 1 in the main paper, namely FA-VAE (Ours), KL-VAE, and VA-VAE, selected based on their overall lowest reconstruction losses. These models represent a diverse set of tokenizers with similar latent spatial compression and achieve top-tier quantitative performance across both reconstruction and perceptual metrics. Without cherry-picking, we present their reconstructions to highlight the perceptual fidelity of FA-VAE compared to the best-performing baselines.
Figure 6: Qualitative reconstructions using KL-VAE on ImageNet $\times256$.
Figure 7: Qualitative reconstructions using VA-VAE on ImageNet $\times256$.
Figure 8: Qualitative reconstructions using our proposed FA-VAE on ImageNet $\times256$.
As shown in Figures 6, 7, and 8, FA-VAE achieves superior visual quality and preserves fine details better than KL-VAE and VA-VAE.
Appendix E Additional Evaluations
To evaluate the reconstruction capabilities and perceptual fidelity of different tokenizer models trained on larger datasets, we report in Table 4 a comparison across frequency-aware metrics (reconstruction, low/high frequency loss) and LPIPS perceptual similarity on the ImageNet (Deng et al. 2009) validation set. Our proposed FA-VAE achieves the best performance across all frequency bands and ranks competitively in perceptual similarity, outperforming prior models such as KL-VAE and VQ-VAE trained on OpenImages (Krasin et al. 2017), as well as DC-AE trained on broader multi-domain datasets (ImageNet (Deng et al. 2009), SAM (Kirillov et al. 2023), FFHQ (Karras, Laine, and Aila 2019), Mapillary Vistas (Neuhold et al. 2017)). These results underscore the effectiveness of frequency-aware latent representations in capturing high perceptual details. We believe our method, if trained on larger datasets, would exhibit even more beneficial results.
The DC-AE, KL-VAE, and VQ-VAE architectures and training setups are adapted from prior work (Xie et al. 2024, 2025; Rombach et al. 2022).